### v2

# Video-XL: Extra-Long Vision Language Model for Hour-Scale Video Understanding

Abstract
Although current Multi-modal Large Language Models (MLLMs) demonstrate promising results in video understanding, processing extremely long videos remains an ongoing challenge. Typically, MLLMs struggle with handling thousands of tokens that exceed the maximum context length of LLMs, and they experience reduced visual clarity due to token aggregation. Another challenge is the high computational cost stemming from the large number of video tokens. To tackle these issues, we propose Video-XL, an extra-long vision language model designed for efficient hour-scale video understanding. Specifically, we argue that LLMs can be adapted as effective visual condensers and introduce Visual Context Latent Summarization, which condenses visual contexts into highly compact forms. Extensive experiments demonstrate that our model achieves promising results on popular long video understanding benchmarks, despite being trained on limited image data. Moreover, Video-XL strikes a promising balance between efficiency and effectiveness, processing 1024 frames on a single 80GB GPU while achieving nearly 100% accuracy in the Needle-in-a-Haystack evaluation. We envision Video-XL becoming a valuable tool for long video applications such as video summarization, surveillance anomaly detection, and Ad placement identification.
Figure 1: Results on the Needle-in-a-haystack evaluation within a single 80G GPU. Compared to other video understanding models, Video-XL can achieve 100% accuracy within 1024 frames.
1 Introduction
In recent years, large language models (LLMs) [1, 2, 3, 4] have demonstrated exceptional capabilities in understanding, reasoning, and generating text across a wide range of open-ended tasks. Building on the broad applicability of LLMs, generative multimodal large language models (MLLMs) [5, 6, 7, 8, 9] have made significant progress in their ability to comprehend and generate content across multiple modalities, particularly images. Unlike static images, videos introduce a temporal dimension, consisting of sequences of frames that are crucial for understanding dynamic visual content in conjunction with textual input. Consequently, numerous studies [10, 11, 12, 13] have sought to adapt MLLMs to address the temporal complexities intrinsic to video sequences.
Most existing models are optimized for short videos, typically under one minute, due to the inherent difficulty in establishing sufficient context for longer videos. To address this challenge, several approaches employ compression strategies, such as reducing the number of visual tokens [14, 15, 16] or introducing specialized memory mechanisms [17, 18] to extend the input visual context length. However, these methods inevitably lead to some degree of information loss during the compression process. Other studies attempt to extend the context length of the language backbone [19] or improve multi-node efficiency during the training process [20]. Nonetheless, they often overlook the computational costs incurred during the inference stage. For instance, LongVA and LongVila can process no more than 500 frames and 200 frames, respectively, on a single 80GB GPU.
Expanding visual context length while maintaining computational efficiency remains a significant challenge. In response, we propose that the LLM itself can serve as an effective visual condenser with appropriate adaptations, leveraging its inherent context representation capabilities. By condensing the LLM’s raw activations (i.e., keys and values from the self-attention module) into highly compact forms, the LLM can perceive information from a broader context, even within a limited context window. While VoCo-LLaMA [21] attempted to distill visual tokens within the LLM, it struggles to handle long videos with thousands of tokens due to the LLM’s limited context window. To address this, we propose Visual Context Latent Summarization, which captures essential information from long visual contexts. Specifically, similar to  [22], we append visual summarization tokens (VST) to the original visual context, condense the raw activations, and process them in a streaming manner using a sliding window with an autoregressive approach. Moreover, to efficiently learn the condensed representation, we avoid introducing additional training strategies and focus solely on refining activation compression during the Supervised Fine-tuning (SFT) stage.
Based on the aboved designs, we propose Video-XL, an extra-long vision language model for efficient hour-scale video understanding. Video-XL has the following highlights:
(i) Video-XL achieves promising performance on long video benchmarks—MLVU [23], Video-MME [24], LongVideoBench [25], and VNBench [26], despite being trained on limited image data.
(ii) Video-XL strikes a promising balance between efficiency and effectiveness, processing 1024 frames on a single 80GB GPU while achieving nearly 100% accuracy in the Needle-in-a-Haystack evaluation.
(iii) Beyond general video understanding tasks, Video-XL shows potential for long video-specific tasks. For instance, it can provide summaries for hour-long movies, detect anomalous clips in surveillance videos, or identify inserted advertisements in long videos.
Figure 2: Overview of Video-XL. Video-XL utilizes a unified visual encoding scheme for single images, multiple images, and videos. Visual tokens generated from the projector are first split into several chunks. The activations from the visual contexts in each chunk are then condensed into Visual Summarization Tokens (VSTs). Conditioned on the accumulated activations between chunks, Video-XL performs next-token prediction in an autoregressive manner.
2 Model
2.1 Architecture and Pipeline
Following the success of the MLLM paradigm of leveraging visual instruction tuning to align vision and language representations, our model contains three main modules: the LLM backbone, Vision Encoder, and Cross-modality Projector.
Large Language Model Backbone. We choose Qwen-2 [27] with 7B size as our LLM as it exhibits strong language understanding and reasoning capabilities.
Vision Encoder. We use CLIP-ViT-L [28] which is universally utilized to encode input images.
Cross-modality Projector. Following LLaVA [7], we leverage a two-layer MLP with a GELU activation function as a cross-modality projector to align the vision encoder and LLM.
To deal with different multimodal data, including single images, multiple images, and videos, we draw inspiration from  [19] to develop a unified encoding scheme. Specifically, for multiple images or videos, we input frames with a resolution of $336\times 336$ directly into CLIP. For single images, we divide them into patches, each with a resolution of $336\times 336$, and then feed these patches into CLIP. Consequently, an N-frame video or N-patch image is encoded into $N\times M$ tokens, where $M$ represents the number of tokens per image. The patch-splitting method follows the approach described in  [19]. For video processing, we sample one frame per second for videos shorter than 128 frames. For longer videos, we uniformly sample 128 frames.
2.2 Visual Context Latent Summarization
Given a visual signal (single image, multi image or video), we can get a sequence of vision tokens $X=\{x_{1},...,x_{n}\}$, where $n$ is the token number generated from projector. The goal is to compress $X$ into fewer yet more compact representations $C$ ($|C|<|X|$). In the following, we elaborate our compression mechanism and learning method to achieve effective, efficient, and flexible compression.
2.2.1 Compression Mechanism
[…]
$[x_{1},\dots,x_{n}]\xrightarrow{\mathrm{Partition}}[X_{1},\dots X_{\lceil n/w\rceil}],~{}X_{i}=[x_{(i-1)w+1},\dots,x_{iw}]=[x^{i}_{1},\dots,x^{i}_{w}].$ (1)
Next, for each chunk $X_{i}$, we determine a compression ratio $\alpha_{i}$ ($w$ is evenly divisible by $\alpha_{i}$). A group of $k_{i}=w/\alpha_{i}$ VSTs, $V_{i}=[\langle\text{vs}\rangle^{i}_{1},\dots,\langle\text{vs}\rangle^{i}_{k_{i}}]$, are inserted into the chunk in an interleaved manner. In other words, one VST is dispatched to the end of every $\alpha_{i}$ vision tokens:
$X_{i}\xrightarrow{\mathrm{Interleave}~{}V_{i}}X_{i}^{\prime}=[x^{i}_{1},\dots,x^{i}_{\alpha_{i}},\langle\text{vs}\rangle^{i}_{1},~{}\dots~{},x^{i}_{w-\alpha_{i}+1},\dots,x^{i}_{w},\langle\text{vs}\rangle^{i}_{k_{i}}].$ (2)
The LLM encodes these chunks one by one. During encoding $X_{i}^{\prime}$, Video-XL reuses all modules of the LLM except employing additional projection matrices ($\boldsymbol{W}_{Q}^{v}$, $\boldsymbol{W}_{K}^{v}$, $\boldsymbol{W}_{V}^{v}$) to transform VST hidden states in the self attention module of each layer. As a result, the information of the visual signal is compressed into VSTs’ activations. After encoding $X_{i}^{\prime}$, we discard activations of all the raw vision tokens ($X_{i}$), while we accumulate the activations of VSTs ($V_{i}$). When encoding the next chunk $X_{i+1}^{\prime}$, the LLM directly conditions on the accumulated VST activations ($V_{\leq i}$) as a proxy to the raw vision tokens $X_{\leq i}$.
2.2.2 Learning Method
Video-XL is learned to optimize the generation quality conditioned on the compressed visual signals. Formally, the next-token prediction is written as:
[…]
$\boldsymbol{\Theta}$ denotes the parameters of the LLM, the vision encoder, the projector, the matrices for VSTs at each layer, $\boldsymbol{W}_{Q}^{v}$, $\boldsymbol{W}_{K}^{v}$, $\boldsymbol{W}_{V}^{v}$, and the token embedding of VST, $\boldsymbol{e}_{\langle\text{vs}\rangle}$ (we use one shared embedding for all VSTs). The standard auto-regression loss is minimized for training. Note that we exclude the beacon tokens from the training loss (setting their labels to -100) because they are only intended for compression. To flexibly support a diverse compression granularities, the compression ratio $\alpha_{i}$ for the $i$-th chunk is randomly sampled from $\{2,4,8\}$ during training. At inference, one can choose one compression ratio according to the specific efficiency requirement and stick to it for all chunks.
3 Training Strategy and Data
We employ a two-stage training strategy consisting of pre-training and fine-tuning. During the pre-training stage, we use the Laion-2M dataset [29] to optimize the projector, where visual embeddings from a pre-trained vision encoder are aligned with text embeddings from the LLM. In the fine-tuning stage, we apply visual instruction tuning to optimize both the projector and all LLM parameters. Additionally, we unfreeze the visual encoder to capture improved visual representations of temporal features. As discussed in Section 2.1, although Video-XL can process various multimodal datasets, we only use image data from Bunny 695k [29] in the training process for efficiency.
Model LLM-size MLVU VideoMME w/o sub. VideoMME w sub.
M-avg G-avg Short Medium Long Avg Short Medium Long Avg
Proprietary Models
GPT-4V [1] - 49.2 5.35 70.5 55.8 53.5 59.9 73.2 59.7 56.9 63.3
GPT-4o [30] - 64.6 5.80 80.0 70.3 65.3 71.9 82.8 76.6 72.1 77.2
Gemini-1.5-Pro [31] - - - 81.7 74.3 67.4 75.0 84.5 81.0 77.4 81.3
Claude3-Opus [32] - 36.5 3.39 71.0 57.4 51.2 60.0 73.5 60.1 54.7 62.9
Open-source MLLMs
VideoChat2 [33] 7B 47.9 3.99 48.3 37.0 33.2 39.5 52.8 39.4 39.2 43.8
LLaMA-VID [34] 7B 33.2 4.22 - - - - - - - -
VideoLLaVA [35] 7B 47.3 3.84 45.3 38.0 36.2 39.9 46.1 40.7 38.1 41.6
Shargpt4Video [36] 7B 46.4 3.77 48.3 36.3 35.0 39.9 53.6 39.3 37.9 43.6
InternVL-Chat-V1.5 [37] 20B 50.4 4.02 60.2 46.4 45.6 47.8 61.7 49.1 46.6 52.4
LongVA [19] 7B 56.3 4.33 61.1 50.4 46.2 52.6 61.6 53.6 47.6 54.3
LongVILA [20] 7B - - 61.8 49.7 39.7 50.5 - - - -
Long-LLaVA [15] 13B - - 61.9 51.4 45.4 52.9 66.2 54.7 50.3 57.1
Video-XL 7B 57.0 4.40 58.2 48.9 45.8 51.0 62.1 51.9 49.3 54.4
Table 1: Experimental results on MLVU and VideoMME benchmarks. For MLVU, we report multi-choice average score (M-avg) and generation score (G-avg) on MLVU-dev. For VideoMME, we report results with and without subtitle assitance.
4 Experimental results
4.1 Benchmarks
We evaluated our methods on several mainstream video understanding benchmarks, including MLVU [23], Video-MME [24], VNBench [26] and LongVideoBench [25].
4.2 Main results
4.2.1 Long Video Benchmark Evaluation
MLVU is a long video understanding benchmark comprising 9 distinct tasks, categorized into Multiple-Choice and Generation tasks. As shown in Table 1, our model achieves leading results in both categories within the 7B model family. Furthermore, our approach surpasses some larger models like Intern-VL [37], showcasing a strong capability in long video understanding.
Video-MME is a comprehensive long video understanding benchmark that encompasses a highly diverse range of video types. It evaluates performance across short, medium, and long video lengths, as well as overall effectiveness. Trained only on image data, Video-XL achieves 51.0% and 54.4% accuracy without and with the assistance of subtitles, achieving competitive results with some state-of-the-art models (See Table 1).
VNBench is a synthetic benchmark designed to evaluate models’ long-context abilities, covering tasks such as retrieval, ordering, and counting. As shown in Table 2, Our model demonstrates leading results among open-source models on ordering and counting tass. Additionally, it surpasses GPT-4V in the final score.
LongVideoBench is a challenging benchmark designed to accurately retrieve and reason over detailed multimodal information from long inputs. With a context input of 256 frames, our model achieves a leading performance among all open-source models on the LongVideoBench-dev task, As demonstrated in Table 2,
Model LLM-size VNBench LongVideoBench
Retrieval Ordering Counting Average
Proprietary Models
Gemini 1.5 Pro [31] - 90.7 72.9 36.7 66.7 64.0
GPT-4o [30] - 95.3 73.4 24.5 64.4 66.7
GPT-4V [1] - 93.7 29.5 23.3 48.9 59.1
Open-source MLLMs
VideoChat2 [33] 7B 32.7 0.4 15.5 20.1 39.3
Video-LLaVA [35] 7B 23.8 1.1 12.4 12.4 39.1
LLaVA-NeXT-Video [38] 7B 44.2 0.4 15.5 20.1 43.5
ST-LLM [39] 7B 51.3 0.0 16.7 22.7 -
LongLLaVA [15] 13B 91.1 35.9 29.2 52.1 -
Video-XL 7B 72.0 46.3 33.5 50.6 48.8
Table 2: Experimental results on VNBench and LongVideoBench.
4.2.2 Longer Context Evaluation
To demonstrate our model’s ability to process extremely long contexts, we conducted a needle-in-the-haystack test to evaluate its performance. As shown in Figure 1, both LLaVA-NexT-Video and LongLLaVA employ a naive position encoding extrapolation algorithm, yet they struggle to comprehend key information when provided with more input context. LongVA fine-tunes the LLM to handle extended input lengths, but it still faces high computational costs, limiting it to processing approximately 400 frames on a single 80G GPU. In contrast, our model achieved 100% accuracy with an 8×8$\times$8 × compression ratio and a 1024-frame input on the same hardware. This suggests that our model strikes an optimal balance between accuracy and computational efficiency.
4.3 Ablation studies
Using the Bunny 695k dataset, we trained two models: one without compression and the other with random compression ratios selected from ${2,4,8}$. For the compressed model, we applied varying compression ratios when testing on the video benchmark MLVU and image benchmarks MME and MMBench. The results can be seen in Table 3. Remarkably, even with an 8×8$\times$8 × compression ratio, the compressed models demonstrated promising performance compared to the baseline.
  Model MLVU MME MMBench
  Baseline 57.0 1534 (395) 71.6
2×2$\times$2 × Com. 56.7 1520 (348) 71.4
4×4$\times$4 × Com. 56.5 1518 (328) 71.4
8×8$\times$8 × Com. 56.0 1501 (320) 70.3
[…]
Table 3: Ablation studies on the impact of Visual Context Latent Summarization.
4.3.1 Visualization Results
We provide some typical cases of long video understanding tasks, including video summarization, surveillance anomaly detection, and Ad placement identification, as shown in Figure 3.
Figure 3: Visualization results of Video-XL.
5 Conclusions
We propose Video-XL, an extra-long vision-language model designed for hour-scale video understanding. By adapting the LLM as a visual condenser, we introduce Visual Context Latent Summarization, which condenses visual contexts into highly compact forms. Despite being trained on limited image data, Video-XL achieves promising results on several long video understanding benchmarks and demonstrates its capability in handling long contexts in the Needle-in-a-Haystack evaluation. In the future, we plan to scale up both the training data and model to further enhance performance in long video understanding.